### v1

# mGTE: Generalized Long-Context Text Representation and Reranking Models for Multilingual Text Retrieval

Abstract
We present systematic efforts in building long-context multilingual text representation model (TRM) and reranker from scratch for text retrieval. We first introduce a text encoder (base size) enhanced with RoPE and unpadding, pre-trained in a native 8192-token context (longer than 512 of previous multilingual encoders). Then we construct a hybrid TRM and a cross-encoder reranker by contrastive learning. Evaluations show that our text encoder outperforms the same-sized previous state-of-the-art XLM-R. Meanwhile, our TRM and reranker match the performance of large-sized state-of-the-art BGE-M3 models and achieve better results on long-context retrieval benchmarks. Further analysis demonstrate that our proposed models exhibit higher efficiency during both training and inference. We believe their efficiency and effectiveness could benefit various researches and industrial applications.
{CJK}
UTF8gbsn
mGTE: Generalized Long-Context Text Representation and Reranking Models for Multilingual Text Retrieval
Xin Zhang1,2, Yanzhao Zhang1, Dingkun Long1, Wen Xie1, Ziqi Dai1, Jialong Tang1, Huan Lin1 Baosong Yang1, Pengjun Xie1, Fei Huang1, Meishan Zhang, Wenjie Li2, Min Zhang 1Alibaba Group, 2Department of Computing, The Hong Kong Polytechnic University {linzhang.zx,zhangyanzhao.zyz,dingkun.ldk}@alibaba-inc.com https://hf.co/Alibaba-NLP/gte-multilingual-base
1 Introduction
Text retrieval aims to find relevant passages or documents from a large corpus given a query Manning (2008). It is often implemented as a multi-stage process, consisting of two main components: a retriever and a reranker Gao et al. (2021a); Zhang et al. (2022); Zhao et al. (2024). The retriever identifies a set of candidate documents that are potentially relevant to the query based on the similarity between their sparse (lexical term weights) or/and dense representations from a text representation model (TRM). While the reranker reorders these retrieved candidates to refine the results based on the relevance score generated by a more precise yet computationally demanding model that processes both the query and a candidate document together.
Recent advances in large language models (LLMs) and retrieval augmented generation (RAG) Gao et al. (2023) systems have led to an unprecedented surge in demand for versatile, plug-and-play TRMs and rerankers. These new applications heavily involve processing long and multilingual texts, which could not be addressed by conventional encoder-based models and urgently require upgraded ones. To this end, some resort to enhancing existing multilingual encoders, e.g., XLM-R Conneau et al. (2020), with extended context window up to 8192 Chen et al. (2024). Others turn to use multilingual LLMs which already have the required capabilities Zhang et al. (2023a), but their models might be computationally expensive for self-hosted search services.
Figure 1: Training pipeline. We first build an 8k long-context multilingual encoder. Then based on it, we train text representation and reranking models for retrieval.
In the English community, it has been proven that training long-context encoders from scratch is promising for text retrieval Günther et al. (2023); Nussbaum et al. (2024). In this work, we continue this journey, presenting systematic efforts in building the long-context multilingual text encoder, TRM, and reranker. We suggest a holistic pipeline (Figure 1) as well as several techniques in modeling and training for multilingual long-context retrieval.
Concretely, we first introduce a text encoder enhanced with RoPE Su et al. (2024) and unpadding Portes et al. (2023), pre-trained by masked language modeling (MLM) Devlin et al. (2019) via a two-stage curriculum for the native 8,192 tokens context. Based on our encoder, we propose a hybrid TRM capable of generating both elastic dense Kusupati et al. (2022) and sparse vectors for efficient first-stage retrieval, as well as a cross-encoder reranker. We construct them via contrastive learning Wang et al. (2022); Li et al. (2023) with large-scale meticulously curated datasets, providing robust off-the-shelf retrieval models.
We conduct extensive experiments to verify our method. For the text encoder, we evaluate on two natural language understanding (NLU) benchmarks, i.e., XTREME-R Ruder et al. (2021) and GLUE Wang et al. (2018), and show that our encoder outperforms the same-sized previous state-of-the-art XLM-R. For the TRM and reranker, we evaluate on multiple retrieval benchmarks with multilingual and long-context settings, e.g., MIRACL Zhang et al. (2023b) and MLDR Chen et al. (2024), where our models match the performance of state-of-the-art BGE-M3 Chen et al. (2024) and achieve better long-context performance by a smaller size. We will open-source our models and code to facilitate further research and applications.
Figure 2: Our text encoder architecture.
2 Method
2.1 Text Encoder
To construct powerful long-context multilingual text encoder models, we implement several enhancements to BERT Devlin et al. (2019) architecture and train it from scratch using the vocabulary of XLM-R111 https://hf.co/FacebookAI/xlm-roberta-base Conneau et al. (2020) series.
Specifically, we replace the absolute positional embeddings with Rotary Position Embedding (RoPE, Su et al., 2024), and upgrade the FFN to gated linear unit (GLU) Shazeer (2020). To ensure compatibility with libraries like FlashAttention (Dao, 2023), we remove the dropout applied to attention scores. In addition, we pad the token embedding size to be a multiple of 64, which could speedup the model throughput Portes et al. (2023).
Unpadding Mode
Inspired by Portes et al. (2023), we unpad the input batch to reduce redundant computations associated with padding tokens (Figure 2). We use xFormers (Lefaudeux et al., 2022) to implement the variable length attention. It dispatch the attention forward and backward to different kernels222 We adopt the memory-efficient attention Rabe and Staats (2021) in this work. based on the numerical precision, attention head size and device type. We unpad the MLM labels as well to reduce the computation cost of predicting non-masked tokens.
Data
We assemble our multilingual pre-training data from a combination of the following sources: C4 (Raffel et al., 2020), Skypile (Wei et al., 2023) (2021-2023 subsets), mC4 (Xue et al., 2021), CulturaX (Nguyen et al., 2023), Wikipedia Foundation and books (proprietary). We filter them and curate a dataset covering 75 Languages. Appendix Table 7 presents the statistics of our dataset.
Training Curriculum
We pre-train the model via masked language modeling (MLM) (Devlin et al., 2019)333 We remove the next sentence prediction objective of BERT following (Liu et al., 2019). . The MLM probability is set to $30\%$ Portes et al. (2023). Following Conneau and Lample (2019) and Conneau et al. (2020), the data from different languages is sampled by a multinomial distribution with probabilities $\{q_{i}\}_{i=1\dots N}$, where
[…]
and $n_{i}$ is the number of texts in language $i$. We set $\alpha=0.5$. This sampling strategy could increase texts from low-resource languages. To train the native 8192-context model more efficiently, we adopt a phased training curriculum Xiong et al. (2024):
  • MLM-2048: we chunk the input into 2048 tokens and set RoPE base to $10,000$.
  • MLM-8192: we chunk the input into 8192 tokens and set RoPE base to $160,000$.
Training Setup
Following Portes et al. (2023), we use the learning rate decoupled444 However, Xie et al. (2024) state that the decoupled weight decay is not ideal. We recommend to keep the default setting. AdamW (Loshchilov and Hutter, 2018) with weight decay $51e-$ 5. We disable gradient clipping (set to 0) (Liu et al., 2019). All models are trained on A100 GPU servers by BF16 PyTorch native automatic mixed precision via transformers Wolf et al. (2020). We list the detailed hyper-parameters of each training stage in Appendix A.2 and Table 8. We denote the resulting models as mGTE-MLM-2048/8192.
Figure 3: Our TRM and reranker.
2.2 Text Representation Model
Based on our encoder, we construct the TRM for the first-stage text retrieval in two steps: contrastive pre-training and fine-tuning Wang et al. (2022); Li et al. (2023). Both steps share the same InfoNCE Oord et al. (2018) learning objective:
$\mathcal{L}=-\log\frac{\text{exp}(s(q,d^{+})/\tau)}{\sum_{i=1}^{N}\text{exp}(s(q,d^{i})/\tau)},$ (2)
where $\tau$, $q$, and $d$ denote the temperature parameter, query and document. The positive $d^{+}$ is the relevant document to $q$, and other irrelevant documents are negatives. These negatives can be either hard-negatives or in-batch negatives (documents of other instances in the same batch). $s(q,d)$ is the relevance score of $q$ and $d$, measured by the dot product or cosine similarity between their respective representations.
Contrastive Pre-Training
We take the encoder output hidden state of the [CLS] token as the dense representation (i.e., embedding) and compute the relevance score by cosine similarity. Our pre-training data (Appendix Table 9) comprise naturally occurring text pairs (e.g., question-answer pairs from Quora and StackExchange, title-content pairs of CommonCrawl), translation pairs Team et al. (2024), and crosslingual instruction tuning data Muennighoff et al. (2022). We train the model with a batch size of $16,384$ and a learning rate of $45e-$ 4 for $240k$ steps. Each batch is sampled from a single data source by the same distribution of Eq.1. The queries (resp. documents) are truncated to the max tokens of $512$ (resp. $1024$). We reverse scale the RoPE base from $160,000$ to $20,000$ to fit the $1024$ context length and acquire the long-context retrieval ability (denotes revNTK, ablation in §3.4). We set $\tau$ of InfoNCE to $0.01$ and only use in-batch negatives. More details refer to Appendix B.3. We denote this contrastive pre-trained model as mGTE-CPT, which is actually an unsupervised embedding model.
Matryoshka Embedding
Many of recently released models and APIs offer elastic embeddings by Matryoshka representation learning (MRL) Kusupati et al. (2022), providing competitive sub-vectors of embeddings to save index storage and speedup search. Let $\bm{e}\in\mathbb{R}^{H}$ denotes an embedding and $\bm{e}_{:d}$ is the sliced sub-vector from dimension 00 to $d<H$. MRL555 Here we mean the MRL-E in Kusupati et al. (2022). optimizes the weighted sum of multiple losses from different $d$ dimensional sub-vectors, i.e., compute InfoNCE by $s_{d}(\bm{e}^{q}_{:d},\bm{e}^{d}_{:d})$. We add this objective to our TRM fine-tuning stage.
Sparse Representation
Chen et al. (2024) show that neural sparse representations (term/token weights predicted by TRM) could greatly improve the long-context retrieval performance. We follow this design, computing the term weight $w_{t}$ of each token of the input by $w_{t}=\text{ReLU}(\bm{W}\bm{h}_{t})$, where $\bm{h}_{t}$ is the encoder hidden state of token $t$ with dimension size $H$ and $\bm{W}\in\mathbb{R}^{H\times 1}$ is randomly initialized. If a token appears multiple times in the text, we keep the max weight. The relevance score is computed by the joint importance of the co-occurring terms (denoted as $q\cap d$) within the query and document pair: $s_{\text{sparse}}(q,d)=\sum_{t\in q\cap d}(w^{q}_{t}\cdot w^{d}_{t})$. This is then used to derive the InfoNCE loss for training.
Contrastive Fine-Tuning
Now we construct the TRM by multi-task learning of matryoshka embedding and sparse representation:
$\mathcal{L}_{\text{TRM}}=\lambda\mathcal{L}_{\text{sparse}}+\sum_{d\in D}w_{d}\mathcal{L}_{:d}\text{~{}},$ (3)
where $D=\{32k\mid k\in\mathbb{N},k\geq 1,32k\leq H\}$ is MRL dimension set, $w_{d}$ is the weight of dimension $d$, and $\lambda$ is the weight of sparse representation loss. We fine-tune our contrastive pre-trained embedding model on diverse high-quality datasets with hard-negatives (e.g., MS MARCO Nguyen et al. (2016), MIRACL Zhang et al. (2023b), listed in Table 11). We adopt a dynamic batching strategy Chen et al. (2024) to fine-tune 8192-context data. The batch sampling strategy is the same as the pre-training stage. The $\tau$ of MRL and sparse is set to $0.05$ and $0.01$ respectively. Other details refer to Appendix B.3. We denote this fine-tuned model as mGTE-TRM.
Model Avg. Pair Class. M.C. Structure Prediction Question Answering Cross-lingual Retrieval
XNLI XCOPA UDPOS WikiANN XQuAD MLQA TyDiQA-GoldP Mewsli-X LAReQA Tatoeba
#Languages (Total 50) 15 11 38 47 11 7 9 38 11 38
Metrics Acc. Acc. F1 F1 F1/EM F1/EM F1/EM mAP@20 mAP@20 Acc.
mBERT-base 59.43 66.63 55.49 71.80 62.34 66.23 / 51.03 57.37 / 42.44 55.01 / 38.05 44.65 75.26 39.49
XLM-R-base 62.02 74.50 50.45 73.84 61.23 72.83 / 58.01 61.54 / 46.45 53.09 / 37.11 42.09 63.43 67.20
mGTE-MLM-2048 65.24 73.17 63.62 73.25 60.87 75.33 / 60.00 64.02 / 48.57 53.58 / 36.68 44.41 72.13 72.02
mGTE-MLM-8192 64.44 73.37 61.98 73.14 59.83 74.81 / 59.37 64.24 / 48.80 49.85 / 33.27 44.52 71.54 71.10
Table 1: XTREME-R Ruder et al. (2021) results in the cross-lingual zero-shot transfer (models are trained on English data) setting. M.C. stands for Multiple Choice. The EM scores are not included in the average.
Model Params Pos. Seq. Len. GLUE Avg.
RoBERTa-baseα 125M Abs. 512 86.4
XLM-R-base 279M Abs. 512 80.44
mGTE-MLM-2048 305M RoPE 2048 83.42
mGTE-MLM-8192 8192 83.47
Table 2: GLUE Wang et al. (2018) devset averages (w/o WNLI). The detailed scores for each subset are shown in Table 13. αTaken from Table 8 of Liu et al. (2019). The rest are from our runs, refer to Appendix C.2.
2.3 Text Reranking Model
We also build a reranker using the cross-encoder architecture. It takes the query and document as input: [CLS] $q$ [SEP] $d$, and directly predicts their relevance score by the [CLS] output state: $s_{\text{rerank}}=\bm{W}\bm{h}_{\text{[CLS]}}$. $\bm{W}\in\mathbb{R}^{H\times 1}$ is randomly initialized. The model is fine-tuned by InfoNCE in one step666 We found that the contrastive pre-training of reranker does not improve the performance. based on our text encoder. Unless otherwise specified, we employ identical data and training settings as the TRM fine-tuning (§2.2). The difference lies in our adjustment of the hard-negatives. We describe the detailed settings in Appendix B.4. We denote this model as mGTE-reranker.
3 Evaluation
We separately evaluate our text encoder in §3.1, TRM and reranker in §3.2 and §3.3.
3.1 Natural Language Understanding
We evaluate the encoder on the cross-lingual natural language understanding (NLU) benchmark XTREME-R777 We use XTREME-R Ruder et al. (2021) instead of XTREME Hu et al. (2020) since we found the retrieval tasks of XTREME is unstable and difficult to evaluate. Ruder et al. (2021) and the English NLU benchmark GLUE Wang et al. (2018). Results show that our encoder outperforms the same-sized previous state-of-the-art XLM-R Conneau et al. (2020) on all benchmarks.
XTREME-R
We focus on the zero-shot cross-lingual transfer setting where models are fine-tuned on English trainset and tested on multi- and cross-lingual data. The fine-tuning setup is described in Appendix C.1. We run mBERT-base, XLM-R-base, and our encoder, as shown in Table 1. Our 2048 and 8192 encoder models achieve average scores that are higher than those of XLM-R by $3.22$ and $2.42$ points, respectively.
GLUE
We also report the performance on the devset of GLUE benchmark Wang et al. (2018). The fine-tuning details refer to Appendix C.2. Table 2 presents the average scores (Table 13 provides the full results). Our models consistently outperform XLM-R-base and reasonably lag behind the English RoBERTa-base Liu et al. (2019).
Model Seq. en zh fr pl
BGE-M3-unsupervised 8192 56.48 57.53 57.95 55.98
mGTE-CPT 512 60.16 58.67 59.72 57.66
8192 60.04 58.63 59.74 57.11
mE5-base 514 59.45 56.21 56.19 55.62
mE5-large 514 61.50 58.81 56.07 60.08
BGE-M3 (Dense) 8192 59.84 60.80 58.79 60.35
mGTE-TRM (Dense) 8192 61.40 62.72 59.79 58.22
E5-mistral-7b 32768 66.63 60.81 48.33 -
voyage-multilingual-2 32000 - - 61.65 -
Cohere-multilingual-v3.0 512 64.01 - 56.02 -
OpenAI-3-large 8191 64.59 - - -
OpenAI-3-small 8191 62.26 - - -
Table 3: Embedding model performance on MTEB English (Muennighoff et al., 2023), Chinese (Xiao et al., 2023), French Ciancone et al. (2024) and Polish (Poświata et al., 2024). The scores of other models are retrieved from the MTEB online leaderboard. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024. Denote our runs.
Params Seq. Len. Avg. MLDR MIRACL MKQA BEIR LoCo
Metric nDCG@10 nDCG@10 recall@20 nDCG@10 nDCG@10
#languages (Total 33) 13 18 25 1 1
BM25 - - 47.0 53.6 31.9 28.1 41.7 79.9
mE5-base 279M 514 53.5 30.5 62.3 53.7 48.9 72.2
mE5-large 560M 514 57.7 34.2 65.4 63.5 51.4 74.3
E5-mistral-7b 7111M 32768 62.4 42.6 62.2 62.4 56.9 87.8
OpenAI-3-large - 8191 - - 54.9 62.1 55.4 79.4
BGE-M3 Dense 568M 8192 64.3 52.5 67.7 67.8 48.7 84.9
BGE-M3 Sparse 55.1 62.2 53.9 36.3 38.3 84.9
BGE-M3 Dense + Sparse 67.7 64.8 68.9 68.1 49.4 87.4
mGTE-TRM Dense 304M 8192 66.7 56.6 62.1 65.8 51.1 88.9
mGTE-TRM Sparse 57.2 71.0 55.9 31.6 39.2 88.1
mGTE-TRM Dense + Sparse 68.9 71.3 64.5 66.0 51.4 91.3
Table 4: Retrieval results on MIRACL Zhang et al. (2023b) and MLDR Chen et al. (2024) (multilingual), MKQA Longpre et al. (2021) (crosslingual), BEIR Thakur et al. (2021) and LoCo Saad-Falcon et al. (2024) (English).
Figure 4: Elastic embedding results on MTEB English.
3.2 Text Embedding
Our contrastive pre-training actually yields a text embedding model. To understand the pre-training and fine-tuning of TRM, and to compare with other models, we first run the most popular text embedding benchmark MTEB Muennighoff et al. (2023) as well as its Chinese, French and Polish versions.
Multilingual MTEB
The results in Table 3 also present the scores of LLM-based models and commercial APIs for reference. For contrastive pre-trained models, our model outperforms BGE-M3-unsupervised Chen et al. (2024) on all four subsets, through our backbone has fewer params than XLM-R-large. Comparing with BGE-M3 and mE5 Wang et al. (2024b), our final TRM achieves best scores on Chinese and French, and is competitive on English.
Elastic Embedding
We compare our TRM (only elastic embeddings) with open-source model and commercial APIs on MTEB English (Figure 4). Our model presents close scores to the same-sized English-only nomic-v1.5, which is promising for a multilingual model. However, it is still behind OpenAI APIs, which is reasonable since they are guessed to be much larger models.
Params Seq. Len. Avg. MLDR MIRACL MKQA BEIR
Metric nDCG@10 nDCG@10 recall@20 nDCG@10
#languages (Total 33) 13 18 25 1
Retrieval (mGTE-TRM Dense) 304M 8192 58.9 56.6 62.1 65.8 50.9
jina-reranker-v2-multilingual 278M 8192 59.4 53.2 65.8 68.8 49.7
bge-reranker-v2-m3 568M 8192 65.7 66.8 72.6 68.7 54.6
mGTE-reranker 304M 8192 67.4 78.7 68.5 67.2 55.4
Table 5: Results of reranking based on the candidates retrieved by our TRM dense model (refer to Table 4).
3.3 Text Retrieval
We conduct evaluations to our TRM and reranker on retrieval benchmarks in multilingual (Miracl Zhang et al. (2023b) and MLDR Chen et al. (2024)), crosslingual (MKQA Longpre et al. (2021)) setting, and the commonly used English BEIR Thakur et al. (2021) and LoCo Saad-Falcon et al. (2024). Our models are close to the state-of-the-art large models on Miracl, MKQA and BEIR, while achieve better scores on long-context datasets MLDR and LoCo. Details are in Appendix E.
First-Stage Retrieval
We compare our TRM to the hybrid model BGE-M3 Chen et al. (2024), dense models like mE5 Wang et al. (2024b) and E5-mistral-7b Wang et al. (2024a), and BM25. As shown in Table 4, our TRM consistently outperforms mE5 and OpenAI APIs, better than BGE-M3 on MLDR, and close to it on the rest parts.
Reranking
In Table 5, we evaluate rerankers based on the candidates retrieved by Our-TRM dense model. Our model outperforms the powerful bge-reranker-v2-m3 Chen et al. (2024) with a smaller size. Moreover, it greatly surpasses the same-sized jina-reranker-v2-multilingual.
Model Attn. Unpad.
Encoding
Time
Search
Latency
BGE-M3 eager $\times$ 1800s 20.35ms
SDPA-MEA 744s
mGTE-TRM eager $\times$ 695s 15.07ms
SDPA-MEA $\times$ 298s
eager 675s
SDPA-MEA 279s
MEA 52s
Table 6: Dense retrieval efficiency. Encoding time is running MLDR-hi corpus (3806 texts with average 4456 tokens after truncating to maximum 8192) on one A100 GPU with FP16. Search latency is measured on a faiss index with 8.8M texts. MEA is the memory-efficient attention in xFormers. SDPA-MEA denotes MEA dispatched by scaled dot-product attention of PyTorch.
3.4 Analysis
Efficiency
We compare the efficiency of our TRM with BGE-M3 on dense retrieval in Table 6. To simulate the real-world scenario, the encoding time is the duration of encoding texts without length grouping. Our TRM is up to 14 times faster than BGE-M3 (52s v.s. 744s). The end-to-end unpadding with xFormers is crucial for encoding, which reduces the time by 5 times (52s v.s. 279s).
Scaled Contrastive Pre-Training
We utilize the reversed NTK scaling in contrastive pre-training to reduce required text length, where we set the RoPE base to $1/8$ of the original and train the 8k encoder with 1k max length. To evaluate the effectiveness, we run the same training without the reversed NTK, comparing the MLDR scores in Figure 5. With revNTK, models exhibit slightly lower performance on 1k context but achieve more stable 8k performance across different training steps.
4 Related Work
Training long-context TRMs has become a hot topic recently. OpenAI released 8191 context APIs Neelakantan et al. (2022) have set the target for open-source community. Portes et al. (2023) and Günther et al. (2023) replace position embedding of BERT with Alibi Press et al. (2022) attention bias and pre-train from scratch, which is shown to be effective in build 8k TRMs. Nussbaum et al. (2024) explore the more powerful RoPE Su et al. (2024) in BERT pre-training and their 2048-context pre-trained encoder achieve better retrieval performance on English. Zhu et al. (2024) suggest patch E5 Wang et al. (2022) with RoPE. We also use RoPE and provide multi-stage training for native 8192-context text encoder, TRM, and reranker.
Chen et al. (2024) propose long-context multilingual TRM and reranker based on XLM-RoBERTa-large Conneau et al. (2020) by extending position embedding to 8192 via continue training. We pre-train native 8k multilingual models from scratch for better long-context performance and efficiency.
Figure 5: MLDR scores in contrastive pre-training. none keeps the RoPE untouched in pre-training. 1024 and 8192 are the max sequence length in evaluations. revNTK-8912 recovers the 8k context by NTK scaling.
5 Conclusion
We present the holistic practice of building native 8192-context multilingual retrieval models. We first suggest a text encoder with RoPE and unpadding, which is pre-trained by a two-stage MLM curriculum for 8k context. Evaluations on NLU benchmarks show that our encoder outperforms XLM-RoBERTa in the same size. Based on our encoder, we construct a hybrid TRM and a cross-encoder reranker by contrastive learning. The TRM is pre-trained with reversed RoPE NTK scaling and fine-tuned to generate both Matryoshka embeddings and sparse representations. Results on monolingual and crosslingual retrieval benchmarks show that our TRM and reranker are close to larger ones on regular datasets, and achieve better performance on long-context datasets. This means our models are more efficient for industrial applications.
Appendix
Appendix A MLM Pre-Training
In this section, we describe the data and training configurations of the MLM pre-training of our suggested text encoder.
A.1 Data
Our multilingual pre-training data are composed from following sources:
  • C4 (Raffel et al., 2020),
  • Skypile (Wei et al., 2023) (2021-2023 subsets),
  • mC4 (Xue et al., 2021) (excluded English),
  • CulturaX (Nguyen et al., 2023),
  • Wikipedia Foundation ,
  • books (proprietary).
We filter them and curate a dataset with 1,028B tokens (by XLM-R tokenizer), covering 75 languages (Chinese Simplified and Traditional are counted as one). Table 7 presents the statistics of our final dataset.
ISO code Language Tokens (M) Size (GiB) ISO code Language Tokens (M) Size (GiB)
af Afrikaans 1,489.19 5.30 ky Kyrgyz 500.40 3.27
ar Arabic 14,549.36 79.53 lo Lao 2.43 0.01
az Azerbaijani 688.72 3.13 lt Lithuanian 1,824.46 6.38
be Belarusian 1,090.61 6.17 lv Latvian 1,823.43 6.38
bg Bulgarian 1,454.57 8.94 mk Macedonian 735.46 4.89
bn Bengali 1,291.58 9.21 ml Malayalam 778.66 7.27
ca Catalan 1,294.05 4.65 mn Mongolian 958.83 5.91
ceb Cebuano 633.06 2.02 mr Marathi 861.05 7.48
cs Czech 1,465.00 5.27 ms Malay 96.37 0.39
cy Welsh 582.49 1.84 my Burmese 902.46 7.26
da Danish 1,030.30 4.01 ne Nepali 657.65 6.32
de German 18,097.31 67.90 nl Dutch 5,137.98 18.65
el Greek 874.87 5.09 no Norwegian 992.51 3.91
en English 187,110.31 771.79 pa Punjabi 726.41 4.96
es Spanish 148,713.06 601.04 pl Polish 2,949.88 10.42
et Estonian 1,111.31 4.10 pt Portuguese 49,594.59 198.64
eu Basque 787.46 2.99 qu Quechua 0.07 0.00
fa Persian 1,203.16 7.22 ro Romanian 2,215.05 7.98
fi Finnish 949.88 3.73 ru Russian 93,966.28 597.92
fr French 136,785.00 512.28 si Sinhala 878.65 7.03
gl Galician 772.47 3.22 sk Slovak 884.38 3.31
gu Gujarati 973.27 6.95 sl Slovenian 1,100.81 4.05
he Hebrew 1,842.74 8.36 so Somali 0.82 0.00
hi Hindi 1,032.67 8.27 sq Albanian 700.78 2.73
hr Croatian 480.19 1.54 sr Serbian 1,139.38 6.84
ht Haitian 0.03 0.00 sv Swedish 840.00 3.37
hu Hungarian 1,341.23 5.10 sw Swahili 31.58 0.13
hy Armenian 805.98 4.88 ta Tamil 926.84 8.54
id Indonesian 25,564.33 119.84 te Telugu 857.91 7.01
is Icelandic 987.89 3.63 th Thai 12,782.08 119.52
it Italian 11,068.23 40.50 tl Filipino 275.16 1.01
ja Japanese 135,684.28 601.19 tr Turkish 1,065.05 4.42
jv Javanese 0.62 0.00 uk Ukrainian 893.70 5.68
ka Georgian 834.90 7.25 ur Urdu 1,051.83 6.19
kk Kazakh 1,020.27 6.57 vi Vietnamese 67,850.87 305.51
km Khmer 746.15 6.54 yo Yoruba 0.04 0.00
kn Kannada 919.83 7.15 zh-cn Chinese (Simplified) 43,727.30 167.23
ko Korean 22,865.85 91.78 zh-tw Chinese (Traditional) 73.39 0.26
Table 7: MLM pre-training data, where we have a total of 1,028B tokens (by XLM-RoBERTa tokenizer). The raw texts are stored in 4.47 TiB arrow files. We report the list of 75 languages (Chinese Simplified and Traditional are counted as one) and include the number of tokens and the size of the data (arrow files, in GiB) for each language.
A.2 Training Details
We pre-train out text encoder with a two-stage curriculum by masked language model (MLM) objective. The first stage model is trained on maximum length $2048$ with batch size $8192$ for roughly $0.6$ epoch (250k steps) on sampled data (by XLM sampling Eq.1). In the second stage, we down sample texts shorter than $2048$ and continue train the model for 30k steps with maximum length $8192$ and batch size $2048$. The RoPE base is set to $10,000$ and $160,000$ for the first and second stage, respectively Xiong et al. (2024); Liu et al. (2024); Men et al. (2024).
The text encoder is initialized in base size ($12$ layers of hidden state size $768$) by PyTorch default initialization. We train the model by transformers library Wolf et al. (2020) in BF16 precision. Following Portes et al. (2023), we use the learning rate decoupled AdamW optimizer with weight decay 1e-5. The other hyper-parameters are in Table 8. During training, we split texts that exceed the max sequence length into chunks, but we do not modify shorter texts.
The 250k steps of first stage, MLM-2048, took 10.75 days on 32 A100 80G GPUs. The 30k steps of second stage, MLM-8192, took 20.5 hours on 32 A100 80G GPUs. We acknowledge that this is not the optimal setting and recommend further explorations to optimize the pre-training.
A.3 Additional Discussion on RoPE
We chose RoPE Su et al. (2024) (to replace absolute position embedding) due to its advantageous properties. RoPE offers excellent context extension capabilities, allowing models to be trained on shorter context windows and then run inference on longer ones. Additionally, it implements asymmetric relative distance encoding, meaning $D(i,j)\neq D(j,i)$, which appears to be particularly important for the training of BERT-like encoder-only models that rely on bidirectional attention. Furthermore, the effectiveness of RoPE has been empirically validated by numerous models, such as RoFormer Su et al. (2024) and LLaMA Touvron et al. (2023).
Hyper-param MLM-2048 MLM-8192
Number of Params 304M
Number of Layers 12
Hidden Size 768
FFN Inner Size 3072
Number of Attention Heads 12
Attention Head Size 64
Dropout 0.1
Attention Dropout 0
Learning Rate Decay Linear
Adam $\epsilon$ 1e-6
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.98
Gradient Clipping 0.0
Precision PyTorch BF16 AMP
Weight Decay 1e-5
Max Length 2048 8192
Batch Size 8192 2048
Peak Learning Rate 5e-4 5e-5
Warm-up Ratio 0.06 0.06
Max Steps 250000 30000
RoPE base 10000 160000
Table 8: MLM pre-training hyper-parameters.
Appendix B Contrastive Learning
In this section, we describe the data and training configurations of the contrastive learning of our TRM and reranker.
B.1 Pre-Training Data
Following previous studies, we create large-scale weakly correlated text pairs from diverse sources. The data are primarily consisted of four parts: English pairs Wang et al. (2022); Li et al. (2023), Chinese pairs Li et al. (2023); Xiao et al. (2023), multilingual pairs (cc-news888 commoncrawl.org/blog/news-dataset-available ), and crosslingual instruction and translation pairs Muennighoff et al. (2022); Team et al. (2024). We filter the data by removing duplicates and low-quality pairs, resulting in a total of 2,938.8M pairs. Table 9 lists the statistics of our contrastive pre-training data (cc-news is separately presented by languages in Table 10).
Source Language Pairs (M) Size (GiB) Source Language Pairs (M) Size (GiB)
agnews English 1.15 0.30 stackoverflow_title_body English 18.01 20.49
amazon_qa English 1.10 0.37 wikihow English 0.13 0.03
amazon_review_title_body English 87.86 43.58 wikipedia English 33.17 19.39
arxiv_title_abstract English 2.26 2.26 yahoo_body_answer English 0.68 0.44
baai_mtp_en English 196.60 178.70 yahoo_qa English 1.20 0.55
beir_dbpedia English 4.64 1.59 yahoo_question_body English 0.66 0.20
beir_debate English 0.38 0.63 baai_mtp_zh Chinese 100.13 231.42
beir_pubmed_title_abstract English 0.13 0.19 baidu_baike Chinese 34.21 39.05
biorxiv_title_abstract English 0.20 0.32 baike_qa_train Chinese 1.43 1.34
clueweb English 3.94 6.62 commoncrawl_zh Chinese 28.42 92.79
clueweb_anchor English 4.51 7.69 gpt3_qa_all Chinese 4.97 2.39
cnn_dailymail English 0.31 1.28 gpt3_summarization Chinese 4.48 1.62
commoncrawl English 139.94 506.84 medical_quac_wenda_10m Chinese 10.00 4.55
dpr_reddit English 199.82 125.71 medical_scholar Chinese 8.43 7.81
gooaq_qa English 3.01 0.97 qcl Chinese 7.40 43.23
hlp_wikipedia English 19.48 13.55 web_text_zh_train Chinese 4.12 2.07
medrxiv_title_abstract English 0.20 0.32 wikipedia Chinese 4.45 1.07
msmarco English 2.89 19.56 wodao Chinese 59.13 190.29
npr English 0.59 1.03 zh_sft_data_v1 Chinese 0.45 0.43
reddit_title_body English 124.89 90.36 zh_sft_data_v2 Chinese 2.24 1.37
s2orc_citation_abstract English 30.58 67.81 zhihu_qa Chinese 53.42 40.99
s2orc_citation_title English 51.03 10.84 zhihu_title_body Chinese 0.94 0.29
s2orc_title_abstract English 41.77 30.29 xp3x Crosslingual 351.87 463.85
stackexchange_qa English 3.00 3.36 translation_eg_NLLB Crosslingual 940.63 323.06
stackexchange_title_body English 4.74 4.00
Table 9: Contrastive pre-training data, where cc-news multilingual data are not included (Table 10). For this Table, we have a total of 2,595.57M pairs (raw texts stored by 2.55 TiB jsonl files).
Lang. Pairs (M) Size (GiB) Lang. Pairs (M) Size (GiB) Lang. Pairs (M) Size (GiB) Lang. Pairs (M) Size (GiB)
ar 20.407 32.45 fy 0.044 0.03 lb 0.048 0.05 sk 1.093 1.16
az 0.401 0.23 gl 0.114 0.20 lt 0.321 0.24 sl 1.046 0.93
be 0.039 0.06 gu 0.061 0.06 lv 0.438 0.37 sq 0.282 0.51
bg 3.005 5.03 he 0.397 0.84 mk 0.173 0.44 sr 0.910 1.09
bn 0.463 0.33 hi 14.253 29.90 ml 0.408 0.48 sv 3.361 2.90
ca 0.909 1.30 hr 1.268 1.77 mr 0.278 0.35 sw 0.059 0.07
cs 1.834 2.18 hu 2.668 3.40 my 0.045 0.04 ta 2.125 1.26
da 1.090 1.58 hy 0.125 0.09 nl 6.700 7.41 te 0.355 0.33
de 39.715 57.98 id 6.048 7.46 nn 0.162 0.12 tg 0.038 0.03
el 7.170 14.93 is 0.100 0.05 no 1.978 2.21 th 0.124 0.17
en 0.615 1.47 it 27.827 40.57 or 0.038 0.03 tl 0.055 0.07
es 55.201 86.87 ja 4.139 3.95 pa 0.036 0.04 tr 23.840 26.81
et 0.950 0.85 ka 0.074 0.06 pl 3.530 5.77 uk 5.021 8.42
eu 0.051 0.02 kn 0.192 0.16 pt 12.611 19.28 ur 1.625 0.87
fa 4.839 7.99 ko 8.605 12.48 ro 6.678 9.15 vi 4.375 7.03
fi 1.532 1.93 ky 0.061 0.03 ru 39.451 65.74 MIX 0.359 0.28
fr 21.242 32.67 la 0.035 0.06 sh 0.220 0.18
Table 10: The cc-news multilingual pairs (343.26M in total, raw texts stored by 512.8 GiB jsonl files), used in contrastive pre-training together with all data of Table 9. MIX denotes the mixed pairs of languages that are less than 1GiB (such as af, ceb). We utilize a very large batch size ($16,384$), and since each batch contains text exclusively from a single source, these low-resource languages might not fill an entire batch. Consequently, we have merged these languages together.
B.2 Fine-Tuning Data
We collect publicly available high-quality dataset as our fine-tune data as detailed in Table 11. For English, we utilize seven datasets:
  • MS MARCO Nguyen et al. (2016),
  • Natural Questions (NQ) Kwiatkowski et al. (2019),
  • TriviaQA Joshi et al. (2017),
  • HotpotQA Yang et al. (2018),
  • SQuAD Rajpurkar et al. (2016),
  • FEVER Thorne et al. (2018),
  • AllNLI from SimCSE Gao et al. (2021b).
For Chinese, we compile six datasets:
Additionally, we incorporate three multilingual datasets: Mr.TyDi Zhang et al. (2021), MIRACL Zhang et al. (2023b), and MLDR Chen et al. (2024). We exclusively use the trainset of each dataset and employ our contrastive pre-trained model to mine hard negatives.
Dataset Language Size
MS MARCO, HotpotQA, NQ, NLI, etc. English 1.4M
DuReader, T2-Ranking, SimCLUE, etc. Chinese 2.0M
MIRACL, Mr.TyDi, MLDR Multilingual 118.9K
Table 11: Specification of training data adopted in Fine-tuning stage.
B.3 TRM Training Setup
Here we separately describe the training setting of the contrastive pre-training and TRM fine-tuning.
Contrastive Pre-Training
In the contrastive pre-training, we train a dense representation model (embedder) which take the [CLS] hidden state as the embedding of the input. We use the same XLM sampling strategy (eq.1) to sample batches from each source of Table 9 or cc-news subset of Table 10, where the texts of one batch only come from one single source, and the batch size is $16,384$. We train the model by transformers with deepspeed ZeRO Rajbhandari et al. (2020) stage 1 in FP16 precision for roughly 0.4 epoch (240k steps, took 154 hours on 16 A100 80G GPUs) of our data (3.93B pairs on sampled data by Eq.1). We use the AdamW optimizer with the learning rate 2e-4, linear decay, and warm-up ratio 0.05. The $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1e-07$. We set gradient clipping to 1.0.
TRM Fine-Tuning
In the fine-tuning stage, we further train our embedding model with high-quality datasets as detailed in §B.2. For each query, we incorporate one positive passage and 8 hard negative passages. To enhance long-context retrieval capabilities and maximize training efficiency, we adopt a dynamic batch size strategy as previous work Chen et al. (2024). Firstly, we group the training data according to their lengths for each dataset. Different batch sizes are then used for varying lengths during training. Additionally, we divide the entire batch into multiple sub-batches, encoding each sub-batch iteratively with gradient checkpointing Chen et al. (2016) and then gather them to get the final batch’s embeddings. We train the embedding model with 10 epochs with 8 A100 80G GPUs. All other hyper-parameters remain consistent with those used in the contrastive pre-training stage. In Table 12, we list the batch size of different length.
B.4 Reranker Training Setup
We utilize the identical fine-tuning dataset for both the reranker and the TRM. For each query, we introduce 10 negative samples, comprising 6 hard negatives and 4 randomly selected negatives. All training parameters expect batch size are kept consistent with those employed for the TRM. The batch sizes are listed in Table 12.
length BS(E) S-BS(R) BS(E) S-BS(R)
0-500 768 256 512 256
500-1000 384 128 384 128
1000-2000 256 64 256 64
2000-3000 160 48 160 48
3000-8000 80 16 80 16
Table 12: Batch size (BS) and sub batch size (S-BS) of different length for embedding (E) and reranker (R) model in the fine-tune stage.
Appendix C NLU Evaluation
We evaluate our text encoder as well as baselines on the multilingual XTREME-R Ruder et al. (2021) and English GLUE Wang et al. (2018) benchmarks. We describe the fine-tuning setup and the evaluation details in the following subsections. The evaluation scripts are available in our github repo101010 github.com/izhx/nlu-evals .
C.1 XTREME-R
We only run XTREME-R Ruder et al. (2021) in the zero-shot cross-lingual transfer learning setting, where models are fine-tuned on English trainset and tested on multi- and cross-lingual data. We compare our encoder with mBERT-base-cased111111 hf.co/google-bert/bert-base-multilingual-cased and XLM-RoBERTa-base121212 hf.co/FacebookAI/xlm-roberta-base . All models are fine-tuned in the same setting and hyper-parameters.
The results are already presented in Table 1.
As XTREME-R has no final release, we implement the evaluation code based on the code of XTREME131313github.com/google-research/xtreme. However, there are some differences in the retrieval evaluation, where our code will deduplicate the retrieval corpus. In addition, we implement the XCOPA in multiple choice, which might be different from XTREME-R. In fine-tuning, if not specified, we use the epoch number of 3, learning rate of 2e-5, batch size of 32, and max sequence length of 128 Hu et al. (2020).
XNLI
We fine-tune the model on MNLI141414 hf.co/datasets/nyu-mll/glue MNLI subset. Williams et al. (2018) trainset and then evaluate the checkpoint on XNLI151515 hf.co/datasets/facebook/xnli Conneau et al. (2018).
XCOPA
We run this data as the multiple choice task. The model is first trained on SIQA161616 hf.co/datasets/allenai/social_i_qa citesap-etal-2019-social and then COPA171717 hf.co/datasets/aps/super_glue copa split. Roemmele et al. (2011) for 5 epochs on each dataset. The checkpoint of COPA is evaluated on XCOPA181818 hf.co/datasets/cambridgeltl/xcopa Ponti et al. (2020).
UDPOS
We extract pos-tagging data from the UD De Marneffe et al. (2021) v2.7 and train the model on trainset of English parts by 10 epochs.
WikiANN
We fine-tune the model on the trainset of English by 10 epochs and evaluate on selected WikiANN Rahimi et al. (2019) testsets191919 hf.co/datasets/unimelb-nlp/wikiann .
XQuAD
We fine-tune on the trainset of SQuAD Rajpurkar et al. (2016) v1.1202020 hf.co/datasets/rajpurkar/squad for 3 epochs with the learning rate 3e-5 and max length 384. Then we evaluate the checkpoint on XQuAD212121 hf.co/datasets/google/xquad Artetxe et al. (2020).
MLQA
We directly evaluate the same checkpoint of XQuAD on MLQA222222 hf.co/datasets/facebook/mlqa Lewis et al. (2020) with the same setting.
TyDiQA-GoldP
We train the model on TyDiQA-GoldP232323 hf.co/datasets/juletxara/tydiqa_xtreme Clark et al. (2020) trainset in the same setting as XQuAD. Then we evaluate the checkpoint on the testset.
Mewsli-X
We generate the data following their github242424 https://github.com/google-research/google-research/blob/master/dense_representations_for_entity_retrieval/mel/mewsli-x.md#getting-started . This is a updated version so that we can not compare with the results in the XTREME-R paper. We train the model on the English wikipedia (mention, entity)-pairs for 2 epochs with the batch size 64 and max length 64. Then we evaluate the checkpoint in the language agnostic retrieval setting, refer to Ruder et al. (2021) for more details.
LAReQA
This task is actually conducted on XQuAD-R252525 hf.co/datasets/google-research-datasets/xquad_r Roy et al. (2020). We fine-tune the model on the trainset of SQuAD v1.1 in dual-encoder architecture ([CLS] as the embedding) and retrieval setting for 3 epochs with the batch size 16, max query length 96, and max document length 256. Then we evaluate the checkpoint on XQuAD-R in same setting.
Tatoeba
We directly evaluate the checkpoint from LAReQA on Tatoeba262626 hf.co/datasets/mteb/tatoeba-bitext-mining Facebook (2019) in the same setting.
C.2 GLUE
The GLUE benchmark Wang et al. (2018) is English transfer learning, i.e., models are trained and tested on the trainset and testset of each dataset (CoLA Warstadt et al. (2019), SST-2 Socher et al. (2013), MRPC Dolan and Brockett (2005), STS-B Cer et al. (2017), QQP, MNLI Williams et al. (2018), QNLI Rajpurkar et al. (2016), RTE).
We evaluate the GLUE benchmark based on the scripts272727 github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification#glue-tasks and data282828 hf.co/datasets/nyu-mll/glue provided by transformers. In fine-tuning of each dataset, we use the epoch number of 3, learning rate of 2e-5, batch size of 32, and max sequence length of 128. For MRPC, STS-B, and RTE, we start from the checkpoint of MNLI following Liu et al. (2019). The MNLI checkpoint is shared with XNLI of XTREME-R (§C.1).
The detailed results are in Table 13. We also include scores of our English models (Our-en-*, pre-trained on C4-en) and baselines Portes et al. (2023); Günther et al. (2023); Nussbaum et al. (2024).
Single Sentence Paraphrase and Similarity Natural Language Inference
Model Params Pos. Seq. Avg. CoLA SST-2 MRPC STS-B QQP MNLI QNLI RTE
RoBERTa-baseα 125M Abs. 512 86.4 63.6 94.8 90.2 91.2 91.9 87.6 92.8 78.7
MosaicBERT-base-128β 137M Alibi 128 85.4 58.2 93.5 89.0 90.3 92.0 85.6 91.4 83.0
MosaicBERT-base-2048γ 137M Alibi 2048 85 54 93 87 90 92 86 92 82
JinaBERT-baseδ 137M Alibi 512 82.6 51.4 94.5 88.4 89.5 80.7 85.7 92.2 78.7
nomic-bert-2048γ 137M RoPE 2048 84 50 93 88 90 92 86 92 82
GTEv1.5-en-base-2048 137M RoPE 2048 85.15 54.46 93.81 93.21 90.00 88.61 86.73 91.67 82.67
GTEv1.5-en-base-8192 137M RoPE 8192 85.61 57.02 93.35 92.14 90.21 88.78 86.69 91.85 84.84
XLM-R-base 279M Abs. 512 80.44 30.74 92.43 92.74 89.16 87.74 84.54 90.37 75.81
mGTE-MLM-2048 305M RoPE 2048 83.42 49.65 92.66 91.17 89.95 88.41 85.40 91.38 78.70
mGTE-MLM-8192 305M RoPE 8192 83.47 48.41 92.32 90.94 89.77 88.50 85.58 91.34 80.87
RoBERTa-largeα 355M Abs. 512 88.9 68.0 96.4 90.9 92.4 92.2 90.2 94.7 86.6
MosaicBERT-large-128β 434M Alibi 128 86.1 59.7 93.7 88.2 90.9 92.0 86.9 93.0 84.5
JinaBERT-largeδ 435M Alibi 512 83.7 59.6 95.0 88.5 88.2 80.9 86.6 92.5 78.5
GTEv1.5-en-large-512 434M RoPE 512 88.16 64.80 94.50 92.09 91.50 89.23 89.12 93.78 90.25
GTEv1.5-en-large-2048 434M RoPE 2048 87.02 60.09 94.61 92.14 91.47 89.12 89.02 92.31 87.36
GTEv1.5-en-large-8192 434M RoPE 8192 87.58 60.39 95.07 93.45 91.37 89.19 89.20 93.90 88.09
Table 13: GLUE Wang et al. (2018) devset scores (w/o WNLI). αTaken from Table 8 of Liu et al. (2019). βTaken from Table S3 of Portes et al. (2023). γTaken from Table 2 of Nussbaum et al. (2024). δTaken from Table 2 of Günther et al. (2023). The rest of the numbers are from our runs, refer to §C.2 for details.
Appendix D Text Embedding Evaluation
We have demonstrated the average scores on MTEB English, Chinese, French and Polish (Table 3). In this section, we delve into the details, presenting results of different tasks on each language. For a fair comparison, we do not include the derived models (developed by secondary training on other public off-the-shelf models) in English and Chinese. In addition to the results obtained from the online leaderboard, our own MTEB evaluations were conducted using version 1.2.0 of mteb library.
MTEB English Param. Dim. Seq. Avg. Class. Clust. PairC. Rerank. Retr. STS Summ.
#Datasets ($\rightarrow$) 56 12 11 3 4 15 10 1
gte-Qwen2-7b-instruct Li et al. (2023) 7613M 3584 131072 70.24 86.58 56.92 85.79 61.42 60.25 83.04 31.35
neural-embedding-v1 - - - 69.94 87.91 54.32 87.68 61.49 58.12 85.24 30.87
NV-Embed-v1 Lee et al. (2024a) 7851M 4096 32768 69.32 87.35 52.8 86.91 60.54 59.36 82.84 31.2
voyage-large-2-instruct - 1024 16000 68.28 81.49 53.35 89.24 60.09 58.28 84.58 30.84
gte-Qwen2-1.5B-instruct Li et al. (2023) 1776M 1536 131072 67.16 82.47 48.75 87.51 59.98 58.29 82.73 31.17
google-gecko Lee et al. (2024b) 1200M 768 2048 66.31 81.17 47.48 87.61 58.9 55.7 85.07 32.63
GritLM-7B Muennighoff et al. (2024) 7242M 4096 32768 66.76 79.46 50.61 87.16 60.49 57.41 83.35 30.37
E5-mistral-7b Wang et al. (2024a) 7111M 4096 32768 66.63 78.47 50.26 88.34 60.21 56.89 84.63 31.4
text-embedding-3-large - 3072 8191 64.59 75.45 49.01 85.72 59.16 55.44 81.73 29.92
mxbai-embed-large-v1 Lee et al. (2024c) 335M 1024 512 64.68 75.64 46.71 87.2 60.11 54.39 85 32.71
nomic-embed-text-v1 Nussbaum et al. (2024) 137M 768 8192 62.39 74.12 43.91 85.15 55.69 52.81 82.06 30.08
gte-en-large-v1.5 434M 1024 8192 65.39 77.75 47.96 84.53 58.5 57.91 81.43 30.91
gte-en-base-v1.5 137M 768 8192 64.11 77.17 46.82 85.33 57.66 54.09 81.97 31.17
mE5-base Wang et al. (2024b) 278M 768 514 59.45 73.02 37.89 83.57 54.84 48.88 80.26 30.11
mE5-large Wang et al. (2024b) 560M 1024 514 61.5 74.81 41.06 84.75 55.86 51.43 81.56 29.69
BGE-m3 (dense) Chen et al. (2024) 568M 1024 8192 59.84 74.08 37.27 84.50 55.28 48.82 81.37 31.55
mGTE-TRM (dense) 305M 768 8192 61.40 70.89 44.31 84.23 57.47 51.08 82.11 30.58
BGE-m3-unsupervised Chen et al. (2024) 560M 1024 8192 56.48 69.28 38.52 80.92 54.03 42.26 78.30 32.11
mGTE-CPT 305M 768 512 60.16 72.89 45.05 84.60 58.41 44.93 80.77 29.94
8192 60.04 72.70 45.35 84.63 58.36 44.46 80.59 30.77
Table 14: Results on MTEB English subset Muennighoff et al. (2023). We compare models from the online leaderboard, where derived models (developed by secondary training on other public off-the-shelf models) are not listed. Denote our runs. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024.
C-MTEB Param. Dim. Seq. Avg. Class. Clust. PairC. Rerank. Retr. STS
#Datasets ($\rightarrow$) 35 9 4 2 4 8 8
gte-Qwen2-7b-instruct Li et al. (2023) 7613M 3584 131072 72.05 75.09 66.06 8 7.48 68.92 76.03 65.33
piccolo-large-zh-v2 Huang et al. (2024) - - - 70.95 74.59 62.17 90.24 70 74.36 63.5
OpenSearch-text-hybrid - 1792 512 68.71 71.74 53.75 88.1 68.27 74.41 62.46
Baichuan-text-embedding - 1024 512 68.34 72.84 56.88 82.32 69.67 73.12 60.07
gte-Qwen2-1.5B-instruct Li et al. (2023) 1776M 1536 131072 67.65 71.12 54.61 86.91 68.21 71.86 60.96
E5-mistral-7b (Wang et al., 2024a) 7111M 4096 32768 60.81 70.17 52.3 72.19 61.86 61.75 50.22
mE5-base Wang et al. (2024b) 278M 768 514 56.21 65.35 40.68 67.07 54.35 61.63 46.49
mE5-large Wang et al. (2024b) 560M 1024 514 58.81 67.34 48.23 69.89 56 63.66 48.29
BGE-m3 (dense) Chen et al. (2024) 568M 1024 8192 60.80 66.95 45.75 73.98 62.88 65.43 52.43
mGTE-TRM (dense) 305M 768 8192 62.72 64.27 47.48 78.34 68.17 71.95 52.73
BGE-m3-unsupervised Chen et al. (2024) 560M 1024 8192 57.53 65.04 47.10 64.09 58.14 61.45 48.42
mGTE-CPT 305M 768 512 58.67 64.64 50.21 63.95 63.77 64.23 46.74
8192 58.63 64.38 49.84 63.99 64.13 64.30 46.77
Table 15: Results on C-MTEB Xiao et al. (2023) (MTEB Chinese). We compare models from the online leaderboard, where derived models (developed by secondary training on other public off-the-shelf models) are not listed. Denote our runs. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024.
F-MTEB Param. Dim. Seq. Avg. Class. Clust. PairC. Rerank. Retr. STS Summ.
#Datasets ($\rightarrow$) 26 6 7 2 2 5 3 1
gte-Qwen2-7b-instruct Li et al. (2023) 7613M 3584 131072 68.25 81.76 55.56 90.43 78.7 55.65 82.31 31.45
gte-Qwen2-1.5B-instruct Li et al. (2023) 1776M 1536 131072 66.6 78.02 55.01 86.88 83.76 52.56 81.26 30.5
voyage-multilingual-2 - 1024 32000 61.65 68.56 46.57 78.66 82.59 54.56 80.13 29.96
voyage-law-2 - 1024 16000 60.58 68.45 44.23 77.3 82.06 52.98 80.29 30.34
mistral-embed - 1024 - 59.41 68.61 44.74 77.32 80.46 46.81 79.56 31.47
E5-mistral-7b (Wang et al., 2024a) 7111M 4096 32768 48.33 57.72 41.16 76.08 62.2 23.44 65.36 32.22
mE5-base Wang et al. (2024b) 278M 768 514 56.19 66.8 42.66 74.82 71.76 41.19 77.22 30.76
mE5-large Wang et al. (2024b) 560M 1024 514 56.07 68.39 38.7 76.19 72.14 42.17 79.37 30.92
BGE-m3 (dense) Chen et al. (2024) 568M 1024 8192 58.79 71.57 36.54 79.78 77.36 51.13 80.78 31.05
mGTE-TRM (dense) 305M 768 8192 59.79 68.72 41.66 79.47 76.47 52.97 81.36 29.74
BGE-m3-unsupervised Chen et al. (2024) 560M 1024 8192 57.95 69.87 38.43 78.51 75.42 50.05 77.18 28.80
mGTE-CPT 305M 768 512 59.72 70.79 41.15 80.29 76.19 53.44 76.87 29.04
8192 59.74 70.69 41.07 79.56 77.10 53.55 77.24 28.74
Table 16: Results on F-MTEB Ciancone et al. (2024) (MTEB French). We compare top-performing models from the online leaderboard. Denote our runs. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024.
MTEB Polish Param. Dim. Seq. Avg. Class. Clust. PairClass. Retr. STS
#Datasets ($\rightarrow$) 26 7 1 4 11 3
gte-Qwen2-7b-instruct Li et al. (2023) 7613M 3584 131072 67.86 77.84 51.36 88.48 54.69 70.86
gte-Qwen2-1.5B-instruct Li et al. (2023) 1776M 1536 131072 64.04 72.29 44.59 84.87 51.88 68.12
mmlw-roberta-large Dadas et al. (2024) 435M 1024 514 63.23 66.39 31.16 89.13 52.71 70.59
mmlw-e5-large Dadas et al. (2024) 560M 1024 514 61.17 61.07 30.62 85.9 52.63 69.98
mmlw-roberta-base Dadas et al. (2024) 124M 768 514 61.05 62.92 33.08 88.14 49.92 70.7
mE5-base Wang et al. (2024b) 278M 768 514 55.62 59.01 24.97 82.15 44.01 65.13
mE5-large Wang et al. (2024b) 560M 1024 514 60.08 63.82 33.88 85.5 48.98 66.91
BGE-m3 (dense) Chen et al. (2024) 568M 1024 8192 60.35 65.15 25.21 86.46 48.51 69.44
mGTE-TRM (dense) 305M 768 8192 58.22 60.15 33.67 85.45 46.40 68.92
BGE-m3-unsupervised Chen et al. (2024) 560M 1024 8192 55.98 60.30 40.17 79.01 43.26 67.05
mGTE-CPT 305M 768 512 57.66 62.72 38.04 79.70 45.55 67.39
8192 57.11 61.55 38.15 79.53 45.29 66.53
Table 17: Results on MTEB Polish subset Poświata et al. (2024) We compare top-performing models from the online leaderboard. Denote our runs. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024.
MTEB-en
Table 14 shows the results on English MTEB Muennighoff et al. (2023). For reference, we include our English embedding models (Our-en-base/large-embed, trained by the two-stage contrastive learning on the English part of our data) and top-performing systems from the online leaderboard. We can see that the multilingual models still have a noticeable gap compared to the English models.
MTEB-zh
Table 15 presents the C-MTEB Xiao et al. (2023) (MTEB Chinese subset) results. We include the results of several LLM-based embedding models and APIs. Given that the Chinese community is also keen on optimizing embedding models, the gap between multilingual models and Chinese models is quite noticeable.
MTEB-fr
Table 16 demonstrates the F-MTEB Ciancone et al. (2024) (MTEB French subset) results. Our TRM dense is comparable to the specialized French API mistral-embed. However, compared to our our-cpt model, the improvement from fine-tuning is not significant.
MTEB-pl
Table 17 lists the Polish MTEB Poświata et al. (2024) results. Our model does not outperform large-sized BGE and mE5. We speculate this may be due to the limited amount of Polish pairs in the contrastive pre-training, resulting in insufficient training.
Appendix E Text Retrieval Evaluation
The retrieval process can be divided into two main stages: recall and reranking. In the recall stage, documents are retrieved using both dense vectors and sparse representations. The final recall score is calculated by weighting the dense retrieval score with a fixed coefficient of 1 and the sparse retrieval score with coefficients ranging from 0.001 to 0.01. Documents not retrieved by either method receive a score of 0. During the ranking stage, the top 100 documents from the recall results are selected as candidates. These candidates are then sorted using our reranker model to produce the final retrieval results.
We present the detail results of MLDR Chen et al. (2024) (multilingual long-context retrieval, Table 18), MKQA Longpre et al. (2021) (multilingual, Table 19), MIRACL Zhang et al. (2023b) (multilingual, Table 20, BEIR Thakur et al. (2021) (English, Table 21) and LoCo Saad-Falcon et al. (2024) (English long-context, Table 22).
Max Length Avg ar de en es fr hi it ja ko pt ru th zh
BM25 8192 53.6 45.1 52.6 57.0 78.0 75.7 43.7 70.9 36.2 25.7 82.6 61.3 33.6 34.6
[…]
BGE-m3-Dense 8192 52.5 47.6 46.1 48.9 74.8 73.8 40.7 62.7 50.9 42.9 74.4 59.5 33.6 26.0
BGE-m3-Sparse 8192 62.2 58.7 53.0 62.1 87.4 82.7 49.6 74.7 53.9 47.9 85.2 72.9 40.3 40.5
BGE-m3-Dense+Sparse 8192 64.8 63.0 56.4 64.2 88.7 84.2 52.3 75.8 58.5 53.1 86.0 75.6 42.9 42.0
mGTE-TRM Dense 8192 56.6 55.0 54.9 51.0 81.2 76.2 45.2 66.7 52.1 46.7 79.1 64.2 35.3 27.4
mGTE-TRM Sparse 8192 71.0 74.3 66.2 66.4 93.6 88.4 61.0 82.2 66.2 64.2 89.9 82.0 47.4 41.8
mGTE-TRM Dense+Sparse 8192 71.3 74.6 66.6 66.5 93.6 88.6 61.6 83.0 66.7 64.6 89.8 82.1 47.7 41.4
+ mGTE-reranker 8192 73.8 76.6 70.4 69.3 96.4 89.6 67.8 81.9 68.1 71.1 90.2 86.1 46.7 44.8
Table 18: Evaluation of multilingual long-doc retrieval on the MLDR Chen et al. (2024) testset (measured by nDCG@10).
Baselines M3-Embedding mGTE-TRM mGTE-reranker
[…]
ar 13.4 33.8 43.8 59.7 44.3 47.6 55.1 61.9 19.5 62.6 61.9 63.0 55.9 17.5 56.0 58.2
da 36.2 55.7 63.3 71.7 63.6 72.3 67.6 71.2 45.1 71.7 71.3 72.0 69.8 37.9 69.7 71.0
de 23.3 53.2 60.2 71.2 62.3 70.8 67.6 69.8 33.2 69.6 70.2 70.4 68.9 27.0 69.1 70.1
es 29.8 55.4 62.3 70.8 63.8 71.6 68.0 69.8 40.3 70.3 70.2 70.7 69.6 35.1 70.0 71.0
fi 33.2 42.8 58.7 67.7 53.0 63.6 65.5 67.8 41.2 68.3 68.4 68.9 64.2 35.3 64.5 64.9
fr 30.3 56.5 62.6 69.5 61.2 72.7 68.2 69.6 43.2 70.1 70.1 70.8 69.8 36.9 70.4 71.0
he 16.1 34.0 50.5 61.4 37.4 32.4 46.3 63.4 24.5 64.4 63.5 64.6 55.4 22.0 55.4 56.5
hu 26.1 46.1 57.1 68.0 55.9 68.3 64.0 67.1 34.5 67.3 67.7 67.9 64.6 28.8 65.0 66.1
it 31.5 53.8 62.0 71.2 61.6 71.3 67.6 69.7 41.5 69.9 69.9 70.3 69.0 36.2 69.2 70.1
ja 14.5 46.3 50.7 63.1 51.7 57.6 64.2 67.0 23.3 67.8 67.1 67.9 65.3 19.5 65.2 67.2
km 20.7 20.6 18.7 18.3 28.2 23.3 25.7 58.5 24.4 59.2 58.9 59.5 53.6 21.9 53.8 54.7
ko 18.3 36.8 44.9 58.9 40.4 49.4 53.9 61.9 24.3 63.2 62.1 63.3 55.9 21.4 56.1 58.9
ms 42.3 53.8 63.7 70.2 62.4 71.1 66.1 71.6 52.5 72.1 71.8 72.3 69.9 47.8 70.2 70.9
nl 42.5 56.9 63.9 73.0 65.0 74.5 68.8 71.3 52.9 71.8 71.7 72.3 70.7 47.4 70.9 71.5
no 38.5 55.2 63.0 71.1 62.0 70.8 67.0 70.7 47.0 71.4 71.1 71.6 69.1 39.7 69.2 70.2
pl 28.7 50.4 60.9 70.5 57.2 71.5 66.1 69.4 36.4 70.0 69.9 70.4 68.4 31.4 68.3 69.6
pt 31.8 52.5 61.0 66.8 58.7 71.6 67.7 69.3 40.2 70.0 69.8 70.6 69.6 34.9 69.6 70.7
ru 21.8 49.8 57.9 70.6 58.7 68.7 65.1 69.4 29.2 70.0 69.4 70.0 68.5 25.8 68.5 69.6
sv 41.1 54.9 62.7 72.0 61.3 73.3 67.8 70.5 49.8 71.3 71.5 71.5 69.5 43.3 69.9 70.6
th 28.4 40.9 54.4 69.7 59.7 57.1 55.2 69.6 34.7 70.5 69.8 70.8 65.0 30.6 65.2 66.9
tr 33.5 45.5 59.9 67.3 59.2 65.5 64.9 68.2 40.9 69.0 69.1 69.6 67.7 36.0 67.7 69.0
vi 33.6 51.3 59.9 68.7 60.0 62.3 63.5 69.6 42.2 70.5 70.2 70.9 69.4 37.6 69.3 70.3
zh_cn 19.4 50.1 55.9 44.3 38.3 61.2 62.7 66.4 26.9 66.7 66.6 67.3 68.2 23.2 68.4 69.5
zh_hk 23.9 50.2 55.5 46.4 38.3 55.9 61.4 65.8 31.2 66.4 65.9 66.7 63.7 27.8 63.8 65.8
zh_tw 22.5 50.6 55.2 45.9 39.0 56.5 61.6 64.8 29.8 65.3 64.9 65.6 63.8 26.6 63.9 65.7
Avg 28.1 47.9 56.3 63.5 53.7 62.4 62.1 67.8 36.3 68.4 68.1 68.8 65.8 31.6 66.0 67.2
Table 19: Recall@20 on MKQA Longpre et al. (2021) dataset for cross-lingual retrieval in all 25 languages. The All of M3-Embedding denotes the hybrid retrieval result of dense, sparse, and multi-vec scores.
Model Avg ar bn en es fa fi fr hi id ja ko ru sw te th zh de yo
BM25 31.9 39.5 48.2 26.7 7.7 28.7 45.8 11.5 35.0 29.7 31.2 37.1 25.6 35.1 38.3 49.1 17.5 12.0 56.1
[…]
OpenAI-3 54.9 - - - - - - - - - - - - - - - - - -
BGE-M3-Dense 67.8 78.4 80.0 56.9 55.5 57.7 78.6 57.8 59.3 56.0 72.8 69.9 70.1 78.6 86.2 82.6 61.7 56.8 60.7
BGE-M3-Sparse 53.9 67.1 68.7 43.7 38.8 45.2 65.3 35.5 48.2 48.9 56.3 61.5 44.5 57.9 79.0 70.9 36.3 32.2 70.0
BGE-M3-Multi-vec 69.0 79.6 81.1 59.4 57.2 58.8 80.1 59.0 61.4 58.2 74.5 71.2 71.2 79.0 87.9 83.0 62.7 57.9 60.4
BGE-M3-Dense+Sparse 68.9 79.6 80.7 58.8 57.5 59.2 79.7 57.6 62.8 58.3 73.9 71.3 69.8 78.5 87.2 83.1 62.5 57.6 61.8
BGE-M3 All 70.0 80.2 81.5 59.8 59.2 60.3 80.4 60.7 63.2 59.1 75.2 72.2 71.7 79.6 88.2 83.8 63.9 59.8 61.5
mGTE-TRM Dense 62.1 71.4 72.7 54.1 51.4 51.2 73.5 53.9 51.6 50.3 65.8 62.7 63.2 69.9 83.0 74.0 60.8 49.7 58.3
mGTE-TRM Sparse 55.9 66.5 70.4 35.6 46.2 40.0 47.6 66.5 39.8 48.9 47.9 59.3 64.3 47.1 59.4 83.0 70.5 73.7 39.9
mGTE-TRM Dense+Sparse 63.5 73.4 75.1 49.9 57.6 62.7 52.0 74.7 53.5 56.4 52.8 67.1 66.7 63.5 69.5 85.2 75.8 58.4 58.8
+ mGTE-reranker 68.5 77.1 63.1 78.6 56.3 72.4 80.3 79.6 58.6 59.1 74.6 75.5 59.4 56.3 56.5 62.2 72.2 86.3 65.1
Table 20: Multi-lingual retrieval performance on the MIRACL Zhang et al. (2023b) dev set (measured by nDCG@10).
BEIR Avg.
Argu-
Ana
Cli-
mate-
Fever
CQA-
Dup-
Stack
DB-
Pedia
Fever FiQA
Hotpot-
QA
MS
MAR-
CO
NF-
Corpus
NQ Quora
Sci-
docs
Sci-
fact
Touche-
2020
Trec-
Covid
gte-Qwen2-7B-instruct 60.25 64.27 45.88 46.43 52.42 95.11 62.03 73.08 45.98 40.6 67 90.09 28.91 79.06 30.57 82.26
NV-Embed-v1 59.36 68.2 34.72 50.51 48.29 87.77 63.1 79.92 46.49 38.04 71.22 89.21 20.19 78.43 28.38 85.88
gte-Qwen2-1.5B-instruct 58.29 69.72 42.91 44.76 48.69 91.57 54.7 68.95 43.36 39.34 64 89.64 24.98 78.44 27.89 85.38
voyage-large-2-instruct 58.28 64.06 32.65 46.6 46.03 91.47 59.76 70.86 40.6 40.32 65.92 87.4 24.32 79.99 39.16 85.07
neural-embedding-v1 58.12 67.21 32.3 49.11 48.05 89.46 58.94 78.87 42 42.6 68.36 89.02 27.69 78.82 24.06 75.33
GritLM-7B 57.41 63.24 30.91 49.42 46.6 82.74 59.95 79.4 41.96 40.89 70.3 89.47 24.41 79.17 27.93 74.8
e5-mistral-7b-instruct 56.89 61.88 38.35 42.97 48.89 87.84 56.59 75.72 43.06 38.62 63.53 89.61 16.3 76.41 26.39 87.25
google-gecko 55.7 62.18 33.21 48.89 47.12 86.96 59.24 71.33 32.58 40.33 61.28 88.18 20.34 75.42 25.86 82.62
text-embedding-3-large 55.44 58.05 30.27 47.54 44.76 87.94 55 71.58 40.24 42.07 61.27 89.05 23.11 77.77 23.35 79.56
gte-en-large-v1.5 57.91 72.11 48.36 42.16 46.3 93.81 63.23 68.18 42.93 36.95 56.08 89.67 26.35 82.43 22.55 77.49
gte-en-base-v1.5 54.09 63.49 40.36 39.52 39.9 94.81 48.65 67.75 42.62 35.88 52.96 88.42 21.92 76.77 25.22 73.13
BM25 41.7 31.5 21.3 29.9 31.3 75.3 23.6 60.3 22.8 32.5 32.9 78.9 15.8 66.5 36.7 65.6
mE5-large 51.43 54.38 25.73 39.68 41.29 82.81 43.8 71.23 43.7 33.99 64.06 88.18 17.47 70.41 23.39 71.33
mE5-base 48.88 44.23 23.86 38.52 40.36 79.44 38.17 68.56 42.27 32.46 60.02 87.65 17.16 69.35 21.35 69.76
BGE-M3 Dense 48.34 53.95 29.52 39.09 39.80 81.38 41.30 69.44 38.32 31.43 60.60 88.57 16.39 64.36 22.63 55.59
BGE-M3 Sparse 38.30 25.08 24.69 27.51 23.21 88.36 26.79 68.45 19.59 27.5 17.98 73.82 8.89 64.37 30.26 48.00
BGE-M3 Dense+Sparse 49.41 53.88 30.21 39.10 39.89 81.24 40.25 70.11 37.62 32.53 59.58 88.62 15.59 65.74 31.12 55.67
mGTE-TRM Dense 51.07 58.36 34.83 38.12 40.11 92.07 44.99 63.03 39.92 36.66 58.10 88.02 18.26 73.42 22.76 57.4
mGTE-TRM Sparse 39.24 40.06 24.17 25.11 20.0 88.32 28.58 64.68 19.39 28.34 19.71 76.84 10.92 67.72 21.52 53.33
mGTE-TRM Dense+Sparse 51.43 58.48 34.89 38.36 39.72 93.14 44.98 65.01 39.99 36.67 56.90 89.05 18.26 73.45 24.09 58.46
+ mGTE-reranker 55.42 58.53 44.93 38.37 45.62 93.9 44.38 74.51 44.99 36.29 65.21 81.67 18.42 75.59 31.29 77.75
BGE-M3-unsupervised 42.26 59.07 23.05 38.10 31.16 59.15 36.57 53.39 27.79 30.67 39.69 86.38 15.08 61.26 17.62 54.90
mGTE-CPT-512,1024 44.93 52.99 17.93 45.01 37.63 34.13 48.38 54.39 31.76 39.01 48.48 86.82 22.95 72.46 18.56 63.46
mGTE-CPT-8192 44.46 55.14 15.85 44.73 38.74 27.42 47.45 55.93 31.79 38.62 49.27 86.81 22.72 73.08 17.08 62.27
Table 21: BEIR benchmark (Thakur et al., 2021) nDCG@10 scores. We include top models from MTEB Retrieval English leaderboard. Denote our runs.
Model Param. Dim. Seq Avg. Tau Scr. Tau Gov. Tau QMS. QASP. Tit. Art. QASP. Abs. Art.
[…]
nomic-embed-text-v1 Nussbaum et al. (2024) 137M 768 8192 85.5 90.9 97.8 44.2 94.9 99.9
text-embedding-3-small - 1536 8192 82.4 92.2 97.7 27.4 95.9 98.9
text-embedding-3-large - 3072 8192 79.4 88.0 93.6 25.5 93.2 96.8
mGTE-en-base-embed 137M 768 8192 87.4 91.8 98.6 49.9 97.1 99.8
mGTE-en-large-embed 434M 1024 8192 86.7 92.6 98.7 44.5 97.8 99.8
[…]
BGE-M3-Dense Chen et al. (2024) 568M 1024 8192 84.9 93.8 97.4 41.9 93.2 98.3
BGE-M3-Sparse Chen et al. (2024) 568M 1024 8192 84.9 95.5 97.9 46.7 85.7 98.9
BGE-M3-Dense+Sparse Chen et al. (2024) 568M 1024 8192 87.4 97.7 98.2 47.7 93.6 99.7
mGTE-TRM Dense 434M 768 8192 88.9 95.1 97.7 58.5 94.6 98.7
mGTE-TRM Sparse 434M 768 8192 88.1 97.6 97.9 60.1 85.5 99.2
mGTE-TRM Dense+Sparse 434M 768 8192 91.3 98.2 98.3 66.5 94.6 98.7
Table 22: The nCDG@10 scores on the LoCo benchmark Saad-Falcon et al. (2024). Denote our runs.